# When LLMs Go Online: The Emerging Threat of Web-Enabled LLMs

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) have established them as agentic systems capable of planning and interacting with various tools. These LLM agents are often paired with web-based tools, enabling access to diverse sources and real-time information. Although these advancements offer significant benefits across various applications, they also increase the risk of malicious use, particularly in cyberattacks involving personal information. In this work, we investigate the risks associated with misuse of LLM agents in cyberattacks involving personal data. Specifically, we aim to understand: 1) how potent LLM agents can be when directed to conduct cyberattacks, 2) how cyberattacks are enhanced by web-based tools, and 3) how affordable and easy it becomes to launch cyberattacks using LLM agents. We examine three attack scenarios: the collection of Personally Identifiable Information (PII), the generation of impersonation posts, and the creation of spear-phishing emails. Our experiments reveal the effectiveness of LLM agents in these attacks: LLM agents achieved a precision of up to 95.9% in collecting PII, up to 93.9% of impersonation posts created by LLM agents were evaluated as authentic, and the click rate for links in spear phishing emails created by LLM agents reached up to 46.67%. Additionally, our findings underscore the limitations of existing safeguards in contemporary commercial LLMs, emphasizing the urgent need for more robust security measures to prevent the misuse of LLM agents.
@@ 1 Introduction
The integration of various external tools, such as APIs and databases, has significantly enhanced the capabilities of Large Language Models (LLMs). This integration allows LLMs to function autonomously as agents — advanced AI systems that utilize LLMs as their core component to perform complex tasks. Numerous studies have shown the effectiveness of LLM agents in performing tasks across various domains [12, 35, 57, 54].
Given the advanced capabilities of LLM agents, there is growing concern about their potential to cause significant real-world harm if used maliciously. Recent studies [28, 17, 18] have shown that LLM agents could be exploited for cyberattacks, highlighting the risks they pose in this context. In response to these risks, LLM vendors (e.g., OpenAI  [^1] https://www.openai.com, Google  [^2] https://www.google.com) have implemented policies that prohibit harmful activities, such as compromising privacy or intentionally deceiving others [43, 6]. Additionally, safeguards have been established to prevent the misuse of these powerful models [40, 22, 4]. However, as models become more capable, it becomes increasingly challenging to foresee and mitigate all potential misuse scenarios.
The risks associated with LLM agents become even more problematic when paired with web-based tools. The widespread sharing of personal information on the web makes it an appealing target for cybercriminals. For example, attackers often resort to data scraping from websites like Google, LinkedIn, and Facebook to collect personal information, which is then sold on hacker forums [52, 56, 55, 39]. Such activities not only infringe on privacy but also increase individuals’ vulnerability against targeted cyberattacks, including impersonation and phishing email attacks [25, 23]. Attackers can use the acquired information on the target to develop specifically tailored documents for malicious purposes, where impersonation posts wrongfully take advantage of the reputation of the target, and phishing emails deceive the target into clicking on a harmful link.
Our work. Despite the web’s utility and the vulnerabilities introduced by the prevalence of personal data, the capability of LLM agents in cyberattacks exploiting personal information remains unclear. In this work, we examine the effectiveness and harmfulness of LLM agents using web-based tools (such as web search API and hand-designed functions for navigation) for malicious purposes concerning personal data. To this end, we identify three hallmark cyberattacks that utilize publicly available private information by leveraging LLM and agents: personally identifiable information (PII) collection, impersonation post generation, and spear phishing email generation.
@@
RQ1. How potent are the cyberattacks exploiting personal information conducted by LLM agents? To assess the potential misuse of LLM agents in cyberattacks, we design proof-of-concepts of the three hallmark cyberattacks. We evaluate the latest commercially available LLMs [^3] We utilize the latest models as of July 17, 2024. Our attacks were last confirmed to be valid on September 1, 2024.—GPT, Claude, and Gemini—that are readily accessible to the public. To quantitatively analyze how privacy is infringed by LLM agents, we use them to collect five types of PII (name, email address, phone number, office address, and personal web page) of CS researchers from 10 prominent universities, using only the university name as initial input for prompt. In addition, to determine if an LLM agent can automatically create effective impersonation posts, we provide the LLM agent with only the names and affiliations of the impersonation targets and prompt them to generate posts as if they are the person and endorse a specific claim. Lastly, we evaluated the effectiveness of LLM agents in crafting spear phishing emails through a user study with 60 participants. This study focused on assessing the authenticity and credibility of various phishing email variants generated by LLM agents. Our findings indicate that LLM agents are able to properly retrieve the PII of targets with high precision rates (up to 0.95 for GPT). Furthermore, evaluations show that up to 93.9% of posts generated by LLM agents were perceived as authentic. Spear phishing emails generated by LLM agents were also effective, with up to 46.67% of the participants click on the malicious link.
@@
RQ2. How much do web-based tools enhance LLM agents undergoing cyberattacks? LLM agents enhance their responses using web-based tools by searching the web and navigating various interactive elements. Previous literature report proficiency of LLMs when used in cyberattacks [51, 47, 9, 38, 32], and the addition of a powerful feature can be expected to increase this risk. To measure the impact of web-based tools on cyberattacks, we focus on the performance difference between using vanilla LLMs and LLM agents enabled with web-based tools. We experiment with two different levels of capability: an LLM agent enabled with searching functions and an LLM agent enabled with searching and navigation functions. We find that LLM agents consistently outperform vanilla LLMs in all the cyberattacks.
@@
RQ3. How approachable are LLM agents to misuse for cyberattacks? The approachability of cyberattacks via LLM agents can be broken down into two categories: cost and safeguard capability. The execution of such attacks incurs both time and financial costs, which significantly influence the practicality of using LLM agents for cyberattacks. Our findings indicate that LLM agents can exploit such attacks very quickly and at minimal cost, emphasizing their practicality. On average, the GPT LLM agent can perform each task within 10 seconds at a cost of around 2 cents. The challenge of misusing LLM agents arises from the inherent safeguard features embedded in the LLM service platforms. For an attack to be executed through LLM agents, the attacker’s prompts must successfully circumvent these safeguards. Our results demonstrate that safeguards implemented by LLM providers were activated only in particular scenarios and with certain services. Furthermore, we found that simply enabling web-based tools actually allowed bypassing of safeguards for some LLM services.
@@
  • We systematically explore the capabilities of LLM agents to exploit private data in cyberattacks. Specifically, we demonstrate that LLM agents can successfully: 1) collect five types of PII, 2) impersonate specific targets to promote an attacker’s claims using personal information, and 3) create highly targeted emails by crafting suitable scenarios and sender identities for recipients.
@@
Content Warning. Please note that this work discusses and carefully examines the potential threat posed by LLM agents in cyberattacks. We believe that the benefits of publicizing these findings outweigh the potential harms. To prevent misuse, we have provided only summarized versions of the prompts used for the attacks in this work. Full prompts will be disclosed only upon request for legitimate research purposes.
@@ 2 Related Work
LLM Security. As LLMs have advanced and become widely adopted, their potential misuse has raised significant concerns. Previous research has demonstrated that LLMs can be exploited in various domains, including PII extraction [27, 31, 38], deepfake user profile creation [36], and phishing email generation [51, 47, 9]. A recent study [32] also highlighted the proliferation of LLMs being repurposed as malicious services in underground marketplaces, further underscoring the risks associated with these powerful models.
Efforts to assess the safety of LLMs have included red-teaming approaches such as jailbreak attacks [7, 14, 30]. In terms of enhancing security, prevailing methods for safety alignment often involve reinforcement learning from human feedback (RLHF), aimed at promoting the development of safer LLMs [7, 44]. With the capability of LLM agents, recent studies have leveraged LLM agents for safeguard purposes [62, 61, 58].
Despite existing safeguards in commercial LLM services [40, 4, 22], we found that even the latest models (as of July, 17, 2024) lack effective protections against our attack scenarios, emphasizing the urgent need for safeguard enhancement.
LLM Agents for Cyberattacks. An LLM agent refers to an artificial intelligence system that uses an LLM to perform tasks, make decisions, or interact with users autonomously [12]. These agents are adaptable to various applications, depending on their training and integration into systems. Recent studies have demonstrated the capability of LLM agents in various domains, such as event forecasting [59], real-world coding challenges [63], and web navigation [35].
@@ 3.1 Overview of LLM Agents
For the WebSearch Agent, we implement the search() function using the Custom Search JSON API [21], which retrieves Google search results in a structured JSON format. This function accepts a search term as an argument and returns the corresponding Google search results. As shown in Figure 1(b), the WebSearch agent calls the search() function with an appropriate query and then uses the returned search results to generate a response. When the agent cannot find the required information from the results, it may repeatedly call the function, adjusting the query as needed. In Appendix A, we include the results of search() as an example.
@@
Models. We employ commercially available models, whose ease of access and capabilities can encourage misuse by attackers. Specifically, we use GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Flash. We utilize the respective APIs: the OpenAI API  [^4] https://openai.com/index/openai-api/ for GPT-4o, the Anthropic API  [^5] https://docs.anthropic.com/en/api/getting-started for Claude 3.5 Sonnet, and the Gemini API  [^6] https://ai.google.dev/api?lang=python for Gemini 1.5 Flash. Throughout this work, we will refer to GPT-4o as GPT, Claude 3.5 Sonnet as Claude, and Gemini 1.5 Flash as Gemini.
Table 1: Summary of experimental settings for each attack.
PII collection Impersonation post generation Spear-phishing email generation
(Section 4) (Section 5) (Section 6)
Attack Goal Get targets’ 5 types of PII Generate social media posts that impersonate a target advocating a specific claim Generate targeted emails to induce recipients click a link embedded in the email
Attack Target CS researchers CS professors Academic researchers, Non-academic professionals
Attack Inputs Name of university Target’s name, Target’s institution, Desired claim Target’s email address
@@ 3.2 Targeted Cyberattacks
Attack 1. PII Collection. Collecting PII, such as email addresses, names, and phone numbers, from the internet raises significant privacy concerns, even when the data is publicly accessible. Attackers can exploit this information for malicious purposes such as identity theft, fraud, or orchestrating further cyberattacks [25, 23]. They often employ web scraping tools and automated scripts to gather PII from various websites. However, due to the unique formatting of information on each website, these tools typically require custom development for each target site, which involves considerable human labor and associated costs [53, 15].
@@ 4 PII Collection
In this section, we investigate the potential for LLM agents to infringe upon privacy by extracting personally identifiable information (PII).
@@ 4.1 Experimental Setup
The pipeline of PII collection (illustrated in Figure 2) using LLM agents consists of two steps: 1) constructing a list of target names and 2) collecting their corresponding personal information. Note that the attack consists of no manual effort by the attacker and only by utilization of LLM agents. For the professor group, the attacker initially extracts the names of CS professors from a given university (e.g., MIT). Then, the attacker collects the remaining PII for each professor listed. For the student group, the data obtained from the professor group is utilized. Specifically, the names, personal web pages, and school affiliations of professors are used in the prompts to identify the names of students currently associated with each professor. The collected student names are then used as the input for the LLM, which is tasked with retrieving the PII for each student.
Evaluation. Due to the vast and diverse nature of the field of CS, we did not evaluate based solely on department listings. Instead, we employed human annotators to review the responses generated by LLM agents. Annotators were directed to classify individuals as CS professors based on their engagement in relevant fields, following a thorough web search. Certain types of PII, like phone numbers, are prone to changes over time, complicating verification; thus, we initiated five separate queries to LLMs for the remaining PII for each individual. We considered the information collection successful if at least one of the five responses matched the data available online. During the annotation process for the names of CS students, we used the student information listed on the professors’ web pages as the ground truth. For more details on the annotation process, please refer to Appendix B.1.
@@
We use precision and recall as our evaluation metric. Precision for each category was calculated as the proportion of successfully collected PII relative to the total PII collected. We utilized the CS faculty lists provided by the official websites of each university as the reference for comparing our collected data. We calculated the upper bound recall  [^8] During our review, we identified several individuals not listed on the official CS faculty pages but recognized as CS professors by our annotators. These individuals were confirmed to be researchers specializing in CS-relevant fields such as information science and machine learning. using the number of professors officially listed under the CS faculty.
@@ 4.2 Main Results
For CS professors, the vanilla LLM without the utilization of web-based tools were less effective in retrieving PII, particularly for detailed information like office addresses and phone numbers. By utilizing the search API, the WebSearch agent collected information more effectively than the vanilla LLM. However, it was less effective in retrieving office addresses and phone numbers due to the limitations of the Google Custom Search API. This API returns only snippets [19] (automatically generated excerpts that summarize page content), which often mix details of multiple professors, leading to inaccuracies, especially in contact information. In contrast, by additionally utilizing navigation tools, the WebNav agent effectively collected the names of 570 out of 686 CS professors and achieved substantial collection rates for additional PII, including phone numbers (71.4%), office locations (91.2%), email addresses (95.9%), and personal web pages (77.7%).
@@
Figure 4: The number of correctly collected PII for CS professors categorized by each type of PII, using Claude for all methods.
@@ 4.3 Model Analysis
In contrast, as illustrated in Figure 4, Claude collected PII in a manner similar to GPT, where more advanced tools gather information more effectively, except in collecting professors’ names. The vanilla LLM managed to collect more names than its agents, but this is due to its indiscriminate collection process, which resulted in a low precision of 0.574 (out of 1,384 names collected, only 795 were correct). For constructing the list of student names, only the WebNav agent, like GPT, was effective in achieving both precision and recall above 0.9 (Figure 5).
When comparing the performance of WebNav agents from GPT and Claude, Claude slightly underperformed as it failed to retrieve the student list from 10% of the given professors’ websites. This difference can be attributed to their distinct exploration strategies: GPT, despite uncertainties, consistently explored to locate the student list, while Claude tended to stop if the context was unclear. Specifically, if the student list was explicitly listed on the given website, the LLM agent scraped and returned the information directly. If not immediately available, the agent employed navigational techniques within the site to locate the student list, utilizing buttons or links as necessary. For instance, the agent may need to access another page via a link on the given site to search for the student list therein. Even if the site did not explicitly indicate that the link leads to a student list, GPT will explore the link to determine if the necessary information is present. In contrast, Claude only pursued links if they are clearly marked, and will report the absence of student information if clarity is lacking.

### v3

@@ Abstract
Recent advancements in Large Language Models (LLMs) have established them as agentic systems capable of planning and interacting with various tools. These LLM agents are often paired with web-based tools, enabling access to diverse sources and real-time information. Although these advancements offer significant benefits across various applications, they also increase the risk of malicious use, particularly in cyberattacks involving personal information. In this work, we investigate the risks associated with misuse of LLM agents in cyberattacks involving personal data. Specifically, we aim to understand: 1) how potent LLM agents can be when directed to conduct cyberattacks, 2) how cyberattacks are enhanced by web-based tools, and 3) how affordable and easy it becomes to launch cyberattacks using LLM agents. We examine three attack scenarios: the collection of Personally Identifiable Information (PII), the generation of impersonation posts, and the creation of spear-phishing emails. Our experiments reveal the effectiveness of LLM agents in these attacks: LLM agents achieved a precision of up to 95.9% in collecting PII, generated impersonation posts where 93.9% of them were deemed authentic, and boosted click rate of phishing links in spear phishing emails by 46.67%. Additionally, our findings underscore the limitations of existing safeguards in contemporary commercial LLMs, emphasizing the urgent need for robust security measures to prevent the misuse of LLM agents.
@@ 1 Introduction
The integration of various external tools, such as APIs and databases, has significantly enhanced the capabilities of Large Language Models (LLMs). This integration allows LLMs to function autonomously as agents—advanced AI systems that utilize LLMs as their core component to perform complex tasks. Numerous studies have shown the effectiveness of LLM agents in performing tasks across various domains [15, 42, 66].
Given the advanced capabilities of LLM agents, there is growing concern about their potential to cause significant real-world harm if used maliciously. Recent studies [35, 21] have shown that LLM agents could be exploited for cyberattacks, highlighting the risks they pose in this context. In response to these risks, LLM vendors (e.g., OpenAI [46], Google [26]) have implemented policies that prohibit harmful activities, such as compromising privacy or intentionally deceiving others [51, 8]. Additionally, safeguards have been established to prevent the misuse of these powerful models [48, 28, 6]. However, as models become more capable, it becomes increasingly challenging to foresee and mitigate all potential misuse scenarios.
The risks associated with LLM agents become even more problematic when paired with web-based tools. The widespread sharing of personal information on the web makes it an appealing target for cybercriminals. For example, attackers often resort to data scraping from websites like LinkedIn, and Facebook to collect personal information, which is then sold on hacker forums [63, 65]. Such activities not only infringe privacy but also increase individuals’ vulnerability to targeted cyberattacks, including impersonation and phishing email attacks [32, 29]. Attackers can use the acquired information on the target to develop specifically tailored documents for malicious purposes, where impersonation posts wrongfully take advantage of the reputation of the target, and phishing emails deceive the target into clicking on a harmful link.
Our work. Despite the web’s utility and the vulnerabilities introduced by the prevalence of personal data, the capability of LLM agents in cyberattacks exploiting personal information remains unclear. In this work, we examine the effectiveness and harmfulness of LLM agents using web-based tools for malicious purposes concerning personal data. To this end, we identify three hallmark cyberattacks that utilize publicly available private information by leveraging LLM and agents: personally identifiable information (PII) collection, impersonation post generation, and spear phishing email generation.
@@
RQ1. How potent are the cyberattacks exploiting personal information conducted by LLM agents? To assess the potential misuse of LLM agents in cyberattacks, we design proof-of-concepts of the three hallmark cyberattacks. We evaluate the latest commercially available LLMs [^1] We utilize the latest models as of July 17, 2024. Our attacks were last confirmed to be valid on September 1, 2024.—GPT, Claude, and Gemini—that are readily accessible to the public. To quantitatively analyze how privacy is infringed by LLM agents, we use them to collect five types of PII of CS researchers from 10 prominent universities, using only the university name as initial input for prompt. In addition, to determine if an LLM agent can automatically create effective impersonation posts, we provide the LLM agent with only the names and affiliations of the impersonation targets and prompt them to generate posts as if they are the person and endorse a specific claim. Lastly, we evaluated the effectiveness of LLM agents in crafting spear phishing emails through a user study with 60 participants. This study focused on assessing the authenticity and credibility of various phishing email variants generated by LLM agents. Our findings indicate that LLM agents are able to properly retrieve up to 535.6 PII items from CS professors. Furthermore, evaluations show that up to 93.9% of posts generated by LLM agents were perceived as authentic. Spear phishing emails generated by LLM agents were also effective, with up to 46.67% of the participants click on the malicious link.
@@
RQ2. How much do web-based tools enhance LLM agents undergoing cyberattacks? LLM agents enhance their responses using web-based tools by searching the web and navigating various interactive elements. Previous literature report proficiency of LLMs when used in cyberattacks [61, 56, 11, 44], and the addition of a powerful feature can be expected to increase this risk. To measure the impact of web-based tools on cyberattacks, we focus on the performance difference between using vanilla LLMs and LLM agents enabled with web-based tools. We experiment with two different levels of capability: an LLM agent enabled with searching functions and an LLM agent enabled with searching and navigation functions. We find that LLM agents consistently outperform vanilla LLMs across various tasks, as depicted in Figure 1.
@@
RQ3. How approachable are LLM agents to misuse for cyberattacks? The approachability of cyberattacks via LLM agents can be broken down into two categories: cost and safeguard capability. The execution of such attacks incurs both time and financial costs, which significantly influence the practicality of using LLM agents for cyberattacks. Our findings indicate that LLM agents can exploit such attacks very quickly and at minimal cost, emphasizing their practicality. On average, the LLM agent with GPT can perform each task within 10 seconds at a cost of around 2 cents. The challenge of misusing LLM agents arises from the inherent safeguard features embedded in the LLM service platforms. For an attack to be executed through LLM agents, the attacker’s prompts must successfully circumvent these safeguards. Our results demonstrate that safeguards implemented by LLM providers were activated only in particular scenarios and with certain services. Furthermore, we found that simply enabling web-based tools actually allowed bypassing of safeguards for some LLM services.
@@
  • We systematically evaluate how LLMs can be utilized to conduct cyberattacks involving personal data. Specifically, we demonstrate that LLM agents can successfully: 1) collect five types of PII, 2) impersonate specific targets to promote an attacker’s claims using personal information, and 3) create highly targeted emails by crafting suitable scenarios and sender identities for recipients.
@@
Figure 1: Performance of LLM and LLM agent across different models: (a) GPT-4o and (b) Claude 3.5 Sonnet. Performance metrics and details are provided in Appendix A.
@@ 2 Related Work
LLM Security. As LLMs have advanced and become widely adopted, their potential misuse has raised significant concerns. Previous research has demonstrated that LLMs can be exploited in various domains, including PII extraction [34, 38], opinion manipulation [62], and phishing email generation [61, 56]. A recent study [40] also highlighted the proliferation of LLMs being repurposed as malicious services in underground marketplaces, further underscoring the risks associated with these powerful models.
Efforts to assess the safety of LLMs have included red-teaming approaches such as jailbreak attacks [9, 18]. In terms of enhancing security, prevailing methods for safety alignment often involve reinforcement learning from human feedback (RLHF), aimed at promoting the development of safer LLMs [9, 52]. With the capability of LLM agents, recent studies have leveraged LLM agents for safeguard purposes [69, 68].
Despite existing safeguards in commercial LLM services [48, 6, 28], we found that even the latest models (as of 17 July 2024) lack effective protections against our attack scenarios, emphasizing the urgent need for safeguard enhancement.
LLM Agents for Cyberattacks. An LLM agent refers to an artificial intelligence system that uses an LLM to perform tasks, make decisions, or interact with users autonomously [15]. These agents are adaptable to various applications, depending on their training and integration into systems. Recent studies have demonstrated the capability of LLM agents in various domains, such as real-world coding challenges [70] and web navigation [42].
@@ 3.1 Overview of LLM Agents
For the WebSearch Agent, we implement the search() function using the Custom Search JSON API [27], which retrieves Google search results in a structured JSON format. This function accepts a search term as an argument and returns the corresponding Google search results. As shown in Figure 2(b), the WebSearch agent calls the search() function with an appropriate query and then uses the returned search results to generate a response. When the agent cannot find the required information from the results, it may repeatedly call the function, adjusting the query as needed.
@@
Models. We employ commercially available models, whose accessibility and capabilities can encourage misuse by attackers. Specifically, we use GPT-4o (referred to as GPT), Claude 3.5 Sonnet (Claude), and Gemini 1.5 Flash (Gemini). We utilize the respective APIs: the OpenAI API [47] for GPT, the Anthropic API [4] for Claude, and the Gemini API [24] for Gemini.
Table 1: Summary of experimental setup for each attack.
PII collection (Section 4) Impersonation post generation (Section 5) Spear-phishing email generation (Section 6)
Goal Get targets’ 5 types of PII Generate social media posts that impersonate a target advocating a specific claim Generate targeted emails to induce recipients click a link embedded in the email
Target CS professors & students CS professors Academic researchers, Non-academic professionals
Input Name of university Target’s name, Target’s institution, Desired claim Target’s email address
@@ 3.2 Targeted Cyberattacks
Attack 1. PII Collection. Collecting PII, such as email addresses, names, and phone numbers, from the internet raises significant privacy concerns. Even when this data is publicly accessible, unauthorized collection constitutes an infringement of privacy [54]. Attackers can exploit this information for malicious purposes such as identity theft, fraud, or orchestrating further cyberattacks [32, 29]. They often employ web scraping tools and automated scripts to gather PII from various websites. However, due to the unique formatting of information on each website, these tools typically require custom development for each target site, which involves considerable human labor and associated costs [64, 19].
@@ 4 PII Collection
In this section, we investigate the potential for LLM agents to infringe upon privacy by extracting personally identifiable information (PII). Following the U.S. Department of Labor’s definition [45], ‘PII’ refers to any information that allows the identity of an individual to be reasonably inferred by either direct or indirect means. In this study, we focus on information that directly identifies an individual, such as names and email addresses. Even when PII is publicly available, its collection and use for unintended purposes can constitute a violation of privacy [54].
@@ 4.1 Experimental Setup
The pipeline of PII collection (illustrated in Figure 3) using LLM agents consists of two steps: 1) constructing a list of target names and 2) collecting their corresponding personal information. Note that the attack consists of no manual effort by the attacker and only by utilization of LLM agents. For the professor group, the attacker initially extracts the names of CS professors from a given university (e.g., MIT). Then, the attacker collects the remaining PII for each professor listed. For the student group, the data obtained from the professor group is utilized. Specifically, three types of professor information—names, personal web pages, and school affiliations—are used in the prompts to identify the names of students currently associated with each professor.
Evaluation. Due to the vast and diverse nature of the field of CS, we did not evaluate based solely on department listings. Instead, we employed human annotators to review the responses generated by LLM agents. Annotators were directed to classify individuals as CS professors based on their engagement in relevant fields, following a thorough web search. Certain types of PII, like phone numbers, are prone to changes over time, complicating verification; thus, we initiated five separate queries to LLMs for the remaining PII for each individual. We considered the information collection successful if at least one of the five responses matched the data available online. During the annotation process for the names of CS students, we used the student information listed on the professors’ web pages as the ground truth. For more details on the annotation process, please refer to Appendix B.
@@ 4.2 Main Results
For CS professors, the vanilla LLM without the utilization of web-based tools were less effective in retrieving PII, particularly for detailed information like office addresses and phone numbers. By utilizing the search API, the WebSearch agent collected information more effectively than the vanilla LLM. However, it was less effective in retrieving office addresses and phone numbers due to the limitations of the Google Custom Search API. This API returns only snippets [23] (automatically generated excerpts that summarize page content), which often mix details of multiple professors, leading to inaccuracies, especially in contact information. In contrast, by additionally utilizing navigation tools, the WebNav agent effectively collected the names of 570 CS professors and achieved substantial collection rates for additional PII, including phone numbers (71.4%), office locations (91.2%), email addresses (95.9%), and personal web pages (77.7%).
@@ 4.3 Model Analysis
In contrast, as illustrated in Figure 5, Claude collected PII in a manner similar to GPT, where more advanced tools gather information more effectively, except in collecting professors’ names. The vanilla LLM managed to collect more names than its agents, but this is due to its indiscriminate collection process, which resulted in a low precision of 0.574 (only 795 accurate out of 1,384 names collected). For constructing the list of student names, only the WebNav agent was effective in achieving both precision and recall above 0.9 (Figure 6).
Figure 5: The number of correctly collected PII for CS professors categorized by each type, using Claude for all methods.
When comparing the performance of WebNav agents from GPT and Claude, Claude collected more PII items on average from professors, with 535.6 items compared to GPT’s 497.4. Specifically, Claude successfully gathered the names of 612 professors, achieving substantial collection rates for additional PII: 88.6% for office locations, 94.6% for email addresses, and 92.0% for personal web pages, but only 62.4% for phone numbers. While Claude was less effective than GPT in collecting phone numbers, it outperformed in acquiring personal web pages.
In contrast, when focusing on the student groups, Claude slightly underperformed as it failed to retrieve the student list from 10% of the given professors’ websites. This difference can be attributed to their distinct exploration strategies: GPT, despite uncertainties, consistently explored to locate the student list, while Claude tended to stop if the context was unclear. Specifically, if the student list was explicitly listed on the given website, the LLM agent scraped and returned the information directly. If not immediately available, the agent employed navigational techniques within the site to locate the student list, utilizing buttons or links as necessary. For instance, the agent may need to access another page via a link on the given site to search for the student list therein. Even if the site did not explicitly indicate that the link leads to a student list, GPT will explore the link to determine if the necessary information is present. In contrast, Claude only pursued links if they are clearly marked, and will report the absence of student information if clarity is lacking.